OpenStudy (anonymous):

Simplify. Cubed root of x^-1y^-2

OpenStudy (jdoe0001):

hmm not sure is simplifyable

OpenStudy (jdoe0001):

\(\Large \bf \sqrt[3]{x^{-1}y^{-2}}\) looks very simplified

OpenStudy (jdoe0001):

got any choices?

OpenStudy (anonymous):

No sir. They might have slipped in a trick question though. Thanks

OpenStudy (jdoe0001):

trick or treat! =)

OpenStudy (jdoe0001):

you can change it... to a rational exponent or say a rational radicand with a rational exponent but that'd not be simplified per se

OpenStudy (jdoe0001):

lemme poke it.... a bit

OpenStudy (jdoe0001):

\(\bf \large { a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\qquad \qquad a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}}\\ \quad \\ \quad \\ \sqrt[3]{x^{-1}y^{-2}}\implies \sqrt[3]{\cfrac{1}{x^1}\cfrac{1}{y^2}}\implies \sqrt[3]{\cfrac{1}{x}} \sqrt[3]{\cfrac{1}{y^2}}\implies \cfrac{\sqrt[3]{1}}{\sqrt[3]{x}}\cdot \cfrac{\sqrt[3]{1}}{\sqrt[3]{y^2}}\\ \quad \\ \cfrac{1}{\sqrt[3]{x}}\cdot \cfrac{1}{\sqrt[3]{y^2}}\implies \cfrac{1}{x^{\frac{1}{3}}}\cdot \cfrac{1}{y^{\frac{2}{3}}}\implies \cfrac{1}{x^{\frac{1}{3}}y^{\frac{2}{3}}}}\)

OpenStudy (jdoe0001):

\(\Large \bf \cfrac{1}{x^{\frac{1}{3}}y^{\frac{2}{3}}}\implies \cfrac{1}{\sqrt[3]{xy^2}}\)

OpenStudy (jdoe0001):

then again... not a simplified version per se